OpenStudy (fanduekisses):

Find the standard form of the equation of the parabola with a focus at (0, -8) and a directrix at y = 8.

OpenStudy (fanduekisses):

i get \[y^{2}=32x\]

OpenStudy (anonymous):

WOP WOP WOP

OpenStudy (sidsiddhartha):

if a parabola have a at focus at (0,-a) and if equation of directrix is ,y=a then standard form of the parabola is \[x^2=-4ay\] so in ur question ,a=8 so equation of the parabola will be \[x^2=-4*8y=-32y\] \[\large x^2=-32y\]

OpenStudy (fanduekisses):

oh soit's vertical

OpenStudy (sidsiddhartha):

it will look like